A tuning fork with a frequency of f = 536 Hz is placed near the top of the tube shown below. The water level is lowered so that the length L slowly increases from an initial value of 20.0 cm. Determine the next two values of L that correspond to resonant modes. (Assume that the speed of sound in air is 343 m/s.) shorter length m longer length m
Answer #1

Lambda = v/f = 343/536 = 0.64 m

For open pipe, L = 63.99 /4 = 16 cm

Shorter length, L = 3 x 64/4 = 48 cm = 0.48 m (Ans)

Next frequency = 5 x 64/4 = 80 cm

Longer length, L = 80 cm = 0.8 m (Ans)

Comment in case any doubt please rate my answer....